The intent of this assignment is just basic accounting and financial statement construction These will get more complicated over the course of the semester, but this one is just to ensure that everyone is on the same page. Pay close attention to the notes at the end of the assignment and the definitions posted to the blackboard site (Ratio Definitions and Case Analysis). You will only need each number on time. Check the syllabus for the due datel Below are numbers from a balance sheet and income statement. Construct the financial ratios requested below. Sales Operating Costs Depreciation Interest Expense Tax Expense Cash Receivables Inventories Fixed Assets, Net Payables Accrued Expenses Long-Term Loan Common Equity 512566 367540 10000 5000 29408 1000 30000 61256 50000 11000 10000 50000 71256Assignment Notes: Answers are correct if they are within 5% of the correct response. most recent submissions to have. Grade appeals will be based on your Excel sheet . * Submit as many times as you would like before the submission deadline. I will take the » Use excel to solve these problem. It reduces rounding problems and is an important skill * Submitted answers should have two decimals places (i.e, 0.01). For things that are percentages, submit them as 2 whole numbers and two decimal places-23.20%. The margin for error on the grade is based off a two-decimal place calculation » Definitions for the ratios are on the course site (towards the bottom). Be careful, ROA, for example has several different definitions in the popular press so stick to these definitions rather than something you find through Google Accruals are debts that are soon due but have not yet been paid out - things like the salary owed to employees between pay periods show up as accruals » Assignment: Prepare an income statement and a balance sheet for is company using the information provided Calculate: 5. Ratio of total debt and liabilities to shareholders equity 6. Interest coverage 7. Net profit margin 8. Sales to total assets (Asset turnover) 9. Return on assets 10. Equity multiplieir 11. Return on equity
Answer #1

Following is the income statement

Sales 512566
Operating costs (367540)
Operating profit (Earning before interest depriciation and tax) 145026
Depriciation (10000)
Earning before interest and tax 135026
Interest exp (5000)
Earning before tax 130026
Tax expense (29408)
Earning after tax or net income $100618

Balance sheet :-

Assets
Current assets
Cash 1000
Receivables 30000
Inventories 61256
Total current assets 92256
Non current assets
Fixed assets, net 50000
Total assets $142256
Liabilities & stockholders equity
Liabilities
Payables 11000
Accrued expenses 10000
Long term loan 50000
Total liabilities 71000
Stockholder's equity
Common equity 71256
Total liabilities & stockholder equity $142256

Following are the ratios :-

5 Ratios of debt to shareholders equity = total liabilities / shareholder's equity

= 71000/71256 = 0.99 or 1

6. Interest coverage = earning before interest and tax / interest expense

= 135026/5000 = 27.01

7. Net profit margin = net income / sales x 100

= 100618/512566x100 = 19.63%.

8. Sales to total assets (asset turnover) = sales / total assets = 512566/142256 = 3.60

9. Return on assets = net earnings / total assets x100

= 100618/142256x100 = 70.73%.

10 . Equity multiplier = total assets / total equity

= 142256/71256 = 1.99 or 2

11 . Return on equity = net income / sales x100

= 100618/71256x100 = 141.21%.